OpenStudy (anonymous):

Use the elimination method to solve the following system of equations. x + 3y – z = 2 x – 2y + 3z = 7 x + 2y – 5z = –21

OpenStudy (anonymous):

well here, x + 3y – z=–21........................1 2 x – 2y + 3z=-21,,,,,,,,,,,,,,,,,,,,,,,2 7 x + 2y – 5z=-21 ..........................3 now multiply eqn 1 by 2 2x+6y-2z=-42.........................4 subtrating 2-4 -8y+5z=21..............................5 multiply eqn 1 by 7 7x+21y-7z=-147........................7 subtrating 3 - 6 -19y+2z=126.............................8 z=(19y+126)/2.........................9 substituting value of z in 5 -8y+{5(19y+126)/2}=21................................... now get value of y substitute the value of y in eqn 5 get value of z substitute value of y and z in 1 get value of x ............................gud luck